A ball is thrown directly downward with an initial speed of 7.95 m/s, from a height of 30.9 m. After what time interval does it strike the ground?

___ s


1
Expert's answer
2022-01-05T04:21:22-0500

Explanations & Calculations


  • The ball falls under gravity which is almost constant for the low heights above earth.
  • Hence you can use any of the 4 motion equations suitably to find the unknown value.
  • You are given the initial velocity, height at which the ball was released, and asked for the time it takes to reach the ground.
  • The hidden quantity/ directly not given but you need to know is the gravitational acceleration "\small g = 9.8 ms^{-2}" .
  • Then you can use "\small s = ut+\frac{1}{2}at^2" downwards from the starting point to find the time. (then the downward direction is +, all the vector quantities direct downward are taken + while any others upwards are taken negative) )

"\qquad\qquad
\begin{aligned}
\small +30.9\,m&=\small +7.95\,ms^{-1}\times t+\frac{1}{2}\times(+9.8)\times t^2\\
\small 0 &=\small 4.9t^2+7.95t-30.9\cdots[\text{a quadratic function}]\\
\small t&=\small \begin{cases}
\small 1.83\,s\\
\small -3.45\,s\cdots[\text{neglected}]
\end{cases}
\end{aligned}"


  • Then the time it would take is "\small 1.83\,s"